OpenStudy (anonymous):

find the equastion in slope intercept form for the line contaning the two points show work (-3,13 ) & (-1,-1)

OpenStudy (anonymous):

slope = (y2 - y1)/(x2 - x1) = (13 - -1)/(-2 - -1) = 14/-1 = -14 so with y = mx + b we have y = -14x + b and using (-1, -1) we have -1 = 14 +b or b = -15 therefore y = -14x - 15

OpenStudy (anonymous):

Okayy can you do this other one for me